RE-ANALYSIS OF THE SPIN-ORBIT PERTURBATION FOR THE PHILLIPS SYSTEM AND THE BALLIK-RAMSAY SYSTEM OF THE SPECTRA OF C_{2}

WANG CHEN, \underline{JIAN TANG} and KENTAROU KAWAGUCHI, Graduate School of Natural Science and Technology Okayama University, 3-1-1 Tsushima-Naka, Okayama 700-8530, Japan.

The Phillips system and the Ballik-Ramsay system of the spectra of C_{2} have been studied extensively before, and the energy difference between the ground X^{1}Σ_{g}^{+} state and the first triplet a^{3}Π_{u} state has been determined by analyzing the spin-orbit interaction between the X^{1}Σ_{g}^{+} and b^{3}Σ_{g}^{−} states. However, the analysis was carried out previously for the individual vibronic bands, and the perturbation parameters < v|H_{SO}|v′ > of the spin-orbit interaction determined for the different vibronic states lead to the very different values of A_{so}^{bX}.^{a}^{,}^{b} In the present study, we re-analyzed the previous spectral data^{a,}^{c}^{,}^{d} by using the overlap integrals (Franck-Condon factors) and r-centroids between the vibronic states of X^{1}Σ_{g}^{+} and b^{3}Σ_{g}^{−} calculated from the RKR potential and by fitting all the vibronic states simultaneously. A new set of molecular parameters was obtained, including the single-valued spin-orbit interaction constant A_{so}^{bX} = 3.067(9) cm^{−1} and the energy difference ΔE = 719.84(6) cm^{−1} between the X^{1}Σ_{g}^{+} and a^{3}Π_{u} states, the latter of which is about 1.5 cm^{−1} larger than the previously determined value.^{a} This new result may guide for searching the forbidden transitions between the singlet and triplet states of C_{2}.

aC. Amiot, J. Chauville and J. -P. Maillard, J. Mol. Spectrosc. \underline{75}, 19 (1979).

bS. P. Davis et al., J. Opt. Sol. Am. B. \underline{5}, 1838 (1988).

cM. Douay, R. Nietmann and P. -F. Bernath, J. Mol. Spectrosc. \underline{131}, 250 (1988).

dM-C. Chan et al., Chem. Phys. Lett. \underline{390}, 340 (2004).